# Meta Pruning via Graph Metanetworks : A Universal Meta Learning Framework for Network Pruning

### v2

@@ B.1 Metanetwork(GNN) Architecture
We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture.
@@ B.2.1 DepGraph and Torch-Pruning
DepGraph (Fang et al., 2023) proposes a general sparsity regularization based structural pruning framework that can be applied to a wide range of neural network architectures, including CNNs, RNNs, GNNs, Transformers, etc. Alongside the paper, the authors released Torch-Pruning, a powerful Python library that enables efficient structural pruning for most modern architectures. In our work, the pruning criterion is designed based on the methodology of DepGraph (Fang et al., 2023), and all pruning operations are implemented using Torch-Pruning.
Figure 7: A picture from (Fang et al., 2023). It shows how we group parameters together in structural pruning.
@@ B.2.2 Structural pruning
In all our experiments, we only do structural pruning. Although we did not conduct experiments with unstructured pruning, our implementations are directly applicable to that setting as well. Structural pruning involves removing parameters in predefined groups, so that the resulting pruned model can be used as a standalone network—without relying on the original model with masks, nor requiring specialized AI accelerators or software to realize reductions in memory footprint and computational cost.
@@
For a more comprehensive theory to find all pruning groups in structural pruning, refer to the DepGraph (Fang et al., 2023) paper (section 3.1 & 3.2).
@@ B.2.6 Pay Attention
Though we draw ideas from Fang et al. (2023) and use their Torch-Pruning libirary, we find their paper is inconsistent with their code. We wrote appendix B.2.3 & B.2.4 strictly based on their code, which is the Torch-Pruning library. If you read the Fang et al. (2023) paper section 3.3, you may find our description a little different from theirs, because their paper is inconsistent with their code, and we are in line with their code.
@@ D.2.5 Hyperparameters
NISP (Yu et al., 2018) 43.2% 1.76
Geometric (He et al., 2019b) 93.59 93.26 -0.33 41.2% 1.70
Polar (Zhuang et al., 2020) 93.80 93.83 0.03 46.8% 1.88
DCP-Adapt (Zhuang et al., 2018) 93.80 93.81 0.01 47.0% 1.89
CP (Li et al., 2017a) 92.80 91.80 -1.00 50.0% 2.00
AMC (He et al., 2018b) 92.80 91.90 -0.90 50.0% 2.00
HRank (Lin et al., 2020a) 93.26 92.17 -1.09 50.0% 2.00
SFP (He et al., 2018a) 93.59 93.36 -0.23 52.6% 2.11
ResRep (Ding et al., 2021) 93.71 93.71 0.00 52.8% 2.12
SCP (Kang & Han, 2020) 93.69 93.23 -0.46 51.5% 2.06
FPGM (He et al., 2019a) 93.59 92.93 -0.66 52.6% 2.11
FPC (He et al., 2020) 93.59 93.24 -0.35 52.9% 2.12
DMC (Gao et al., 2020) 93.62 92.69 -0.93 50.0% 2.00
GNN-RL (Yu et al., 2022) 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL (Fang et al., 2023) 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL (Fang et al., 2023) 93.53 93.77 0.24 52.6% 2.11
ATO (Wu et al., 2024) 93.50 93.74 0.24 55.0% 2.22
@@
GBN (You et al., 2019a) 93.10 92.77 -0.33 60.2% 2.51
AFP (Ding et al., 2018) 93.93 92.94 -0.99 60.9% 2.56
C-SGD (Ding et al., 2019) 93.39 93.44 0.05 60.8% 2.55
Greg-1 (Wang et al., 2021) 93.36 93.18 -0.18 60.8% 2.55
Greg-2 (Wang et al., 2021) 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL (Fang et al., 2023) 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL (Fang et al., 2023) 93.53 93.64 0.11 61.1% 2.57
ATO (Wu et al., 2024) 93.50 93.48 -0.02 65.3% 2.88
@@ D.3.5 Hyperparameters
OBD (Wang et al., 2019) 73.34 60.70 -12.64 82.55% 5.73
OBD (Wang et al., 2019) 73.34 60.66 -12.68 83.58% 6.09
EigenD (Wang et al., 2019) 73.34 65.18 -8.16 88.64% 8.80
Greg-1 (Wang et al., 2021) 74.02 67.55 -6.67 88.69% 8.84
Greg-2 (Wang et al., 2021) 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL (Fang et al., 2023) 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL (Fang et al., 2023) 73.50 70.39 -3.11 88.79% 8.92
@@ D.4.5 Hyperparameters
DCP (Zhuang et al., 2018) 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP (Peng et al., 2019) 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM (He et al., 2019a) 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP (Lin et al., 2020b) 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC (Gao et al., 2020) 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random (Li et al., 2022) 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph (Fang et al., 2023) 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO (Wu et al., 2024) 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP (Li et al., 2023) 76.13%(-) 75.55%(-0.58%) - 56.7%
@@
Final Pruning metanetwork epoch 12
@@ G.5 Full Results
All results are summarized in Table 17. Note that unlike CNNs, Vision Transformers naturally lacks inductive bias and need much more epochs of finetuning to reach higher accuracy. For other works like DepGraph (Fang et al., 2023), theirs method need to finetune 600 epochs in total, among them, 300 epochs are sparsity learning (slower than usual finetuning) and it takes more than a week to run a single round. We don’t have enough time to fully run the experiments with more finetuning for now, but we will continue tune our hyperparameters, add more epochs and hopefully give a better results in the next version of our paper.
@@ G.6 Hyperparameters
ViT-B/16 (Dosovitskiy et al., 2021) 81.07% - - 17.6
CP-ViT (Song et al., 2022) 77.91% 77.36% -0.55% 11.7
DepGraph (Fang et al., 2023) 81.07% 79.17% -1.90% 10.4
MetaPruning(ours) 81.07% 77.59% -2.83% 11.0
@@
milestone "3"
pruner reg 1000
Final Pruning metanetwork epoch 6
speed up 1.6000
Finetuning After Metanetwork epoch 100
lr 0.0001
@@
Appendix H The Use of Large Language Models

### v3

@@ B.1 Metanetwork(GNN) Architecture
We build our metanetwork based on the message passing framework (MessagePassing) and PNA (PNA) architecture.
@@ B.2.1 DepGraph and Torch-Pruning
DepGraph (DepGraph) proposes a general sparsity regularization based structural pruning framework that can be applied to a wide range of neural network architectures, including CNNs, RNNs, GNNs, Transformers, etc. Alongside the paper, the authors released Torch-Pruning, a powerful Python library that enables efficient structural pruning for most modern architectures. In our work, the pruning criterion is designed based on the methodology of DepGraph (DepGraph), and all pruning operations are implemented using Torch-Pruning.
Figure 9: A picture from (DepGraph). It shows how we group parameters together in structural pruning.
@@ B.2.2 Structural pruning
We perform most of our experiments by default using structural pruning, but also include a small number of experiments with unstructured and N:M sparsity pruning to demonstrate that our method can be applied to all kinds of pruning. Here we mainly introduce the structural pruning.
Structural pruning involves removing parameters in predefined groups, so that the resulting pruned model can be used as a standalone network—without relying on the original model with masks, nor requiring specialized AI accelerators or software to realize reductions in memory footprint and computational cost.
@@
For a more comprehensive theory to find all pruning groups in structural pruning, refer to the DepGraph (DepGraph) paper (section 3.1 & 3.2).
@@ B.2.6 Pay Attention
Though we draw ideas from DepGraph and use their Torch-Pruning libirary, we find their paper is slightly inconsistent with their code. We wrote appendix B.2.3 & B.2.4 strictly based on their code, which is the Torch-Pruning library. If you read the DepGraph paper section 3.3, you may find our description a little different from theirs, because their paper is slightly inconsistent with their code, and we are in line with their code. DepGraph is really a good paper. The small inconsistency may are made unintentionally by the author, not on purpose.
@@ D.1.1 General Settings
Relationship between "Acc VS. Speed Up Curve" and Pruning Performance: Empirically, pruning performance can be qualitatively predicted from the "Accuracy VS. Speed Up" curve. The accuracy in the flat region of the curve typically represents the maximum achievable accuracy; after pruning and finetuning, the final accuracy is usually the same or slightly lower, and sometimes only a little bit higher at most. The amount of speed up that preserves this top accuracy is generally larger than the speed up at the curve’s turning point, and these two values tend to correlate—the larger the turning point speed up, the larger the maintainable speed up. Therefore, if we aim for a pruned model with, for example, 93% accuracy, we should select a metanetwork whose flat-region accuracy is around or slightly above 93%, and whose flat region is as long as possible. We cannot give a teoretical guarantee between the "Acc VS. Speed Up" curve and the pruning performance. But this is not only our problem, because as far as we know, all other pruning methods can’t give theoretical gurarantee between their pruning criterion and final performance as well. The reasons can be the network itself is too complex and processes like finetuning changes everything in an unpredictable way.
@@ D.2.5 Hyperparameters
NISP (NISP) 43.2% 1.76
Geometric (Geometric) 93.59 93.26 -0.33 41.2% 1.70
Polar (Polar) 93.80 93.83 0.03 46.8% 1.88
DCP-Adapt (DCP-Adapt) 93.80 93.81 0.01 47.0% 1.89
CP (CP) 92.80 91.80 -1.00 50.0% 2.00
AMC (AMC) 92.80 91.90 -0.90 50.0% 2.00
HRank (HRank) 93.26 92.17 -1.09 50.0% 2.00
SFP (SFP) 93.59 93.36 -0.23 52.6% 2.11
ResRep (ResRep) 93.71 93.71 0.00 52.8% 2.12
SCP (SCP) 93.69 93.23 -0.46 51.5% 2.06
FPGM (FPGM) 93.59 92.93 -0.66 52.6% 2.11
FPC (FPC) 93.59 93.24 -0.35 52.9% 2.12
DMC (DMC) 93.62 92.69 -0.93 50.0% 2.00
GNN-RL (GNN-RL) 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL (DepGraph) 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL (DepGraph) 93.53 93.77 0.24 52.6% 2.11
ATO (ATO) 93.50 93.74 0.24 55.0% 2.22
@@
GBN (GBN) 93.10 92.77 -0.33 60.2% 2.51
AFP (AFP) 93.93 92.94 -0.99 60.9% 2.56
C-SGD (C-SGD) 93.39 93.44 0.05 60.8% 2.55
Greg-1 (Greg) 93.36 93.18 -0.18 60.8% 2.55
Greg-2 (Greg) 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL (DepGraph) 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL (DepGraph) 93.53 93.64 0.11 61.1% 2.57
ATO (ATO) 93.50 93.48 -0.02 65.3% 2.88
@@ D.3.5 Hyperparameters
OBD (OBD) 73.34 60.70 -12.64 82.55% 5.73
OBD (OBD) 73.34 60.66 -12.68 83.58% 6.09
EigenD (OBD) 73.34 65.18 -8.16 88.64% 8.80
Greg-1 (Greg) 74.02 67.55 -6.67 88.69% 8.84
Greg-2 (Greg) 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL (DepGraph) 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL (DepGraph) 73.50 70.39 -3.11 88.79% 8.92
@@ D.4.5 Hyperparameters
DCP (DCP-Adapt) 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP (CCP) 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM (FPGM) 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP (ABCP) 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC (DMC) 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random (Random) 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph (DepGraph) 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO (ATO) 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP (DTP) 76.13%(-) 75.55%(-0.58%) - 56.7%
@@
Final Pruning metanetwork epoch 16
@@ F.2 Transfer between Architectures
F.3 Transfer From Small Dataset to Large Dataset
When we are using large dataset like imagenet, is it possible that we use a much smaller dataset during meta-training but get the same results? Our answer is yes.
When pruning Resnet50 on IMAGENET, we evenly choose 10% on each class in IMAGENET and form a new subset dataset. Meta-Training using this subset causes almost no drop in accuracy. See results in Table 19. This suggest that even 10% percent dataset is enough in meta-train to let metanetwork learn the pruning strategy.
Table 19: During meta-train, use full ImageNet vs. only 10% ImageNet
ours 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
ours(10%) 76.14%(93.11%) 76.24%(+0.10%) 92.65%(-0.46%) 57.0%
Table 20: Transfer classification to detection
Method Origin Prune w/o metanetwork Prune with metanetwork
mAP 0.6061 0.4524 0.5173
@@ G.5 Full Results
All results are summarized in Table 21.
@@ G.6 Hyperparameters
ViT-B/16 (ViT) 81.07% - - 17.6
CP-ViT (Cp-vit) 77.91% 77.36% -0.55% 11.7
DepGraph (DepGraph) 81.07% 79.17% -1.90% 10.4
MetaPruning(ours) 81.07% 78.26% -2.81% 10.3
@@
milestone -
pruner reg 10
Final Pruning metanetwork epoch 22
speed up 1.7
Finetuning After Metanetwork epoch 300
lr 0.01
@@
label smoothing 0.1
mixup alpha 0.2
cutmix alpha 0.1
@@
label smoothing 0.1
mixup alpha 0.2
cutmix alpha 0.1
Appendix H More Visualization of Statistics
Figure 13: $l_{1}$ Norm
Figure 14: Inter-Channel Correlation
Figure 15: Effeicient Rank
Appendix I The Use of Large Language Models